OpenStudy (anonymous):

Julie drives 100 mi to Grandmother's house. On the way to Grandmother's, Julie drives half the distance at 33.0 and half the distance at 80.0 . On her return trip, she drives half the time at 33.0 and half the time at 80.0 . What is Julie's average speed on the way to Grandmother's house?

OpenStudy (turingtest):

reprint already answered

OpenStudy (turingtest):

...unless someone wants to plug in the numbers for you that is http://openstudy.com/study#/updates/4f1e3b56e4b04992dd24a260

OpenStudy (anonymous):

(33.0+80.0)/2=56.5 from this formula ==> average speed = (V1+V2+V3+...)/n

OpenStudy (jamesj):

Very importantly, what alismart has written is false in general. For example, if I drive - 30 km in one hour (v1 = 30 km/hr) followed by - 30 km in the next 1/2 hour (v2 = 60 km/hr) then my average speed from first principles is Average speed = (distance)/(time) = (60 km)/(1.5 hours) = 40 km/hr Using alismart's formula, the average speed would be \[ \frac{30 + 60}{2} = 45 \] which is wrong.

OpenStudy (anonymous):

then if you want to bring distance the complete formula will be : x=distance ==>

OpenStudy (anonymous):